he accounting records for Portland Products report the following manufacturing costs for the past year.

Direct materials $ 320,000
Direct labor 260,000
Variable overhead 239,000

Production was 190,000 units. Fixed manufacturing overhead was $724,000.

For the coming year, costs are expected to increase as follows: direct materials costs by 20 percent, excluding any effect of volume changes; direct labor by 4 percent; and fixed manufacturing overhead by 10 percent. Variable manufacturing overhead per unit is expected to remain the same.


Required:

a. Prepare a cost estimate for a volume level of 152,000 units of product this year.
b. Determine the costs per unit for last year and for this year.
Answer #1

At Production- 190,000 Units

Material cost per unit = 320000/190000 = 1.68

Direct labour cost per unit = 260000/190000 = 1.37

Variable overhead unit = 239000/190000 = 1.26

Fixed cost per unit = 724000/190000 = 3.81

Material cost increment 20% = 1.68 + 20% = 2.02

Labour cost increment 4% = 1.37 + 4% = 1.42

Fixed overhead increment 10% = 724000 + 10% = 796400

Fixed overhead per unit = 796400/152000 = 5.24

At Volume level 152,000 Units.

Total cost per unit =$ 2.02 + 1.42 + 1.26 = $4.70

152000 * $4.70= $714,400

Total cost estimate = $714,400 + $796,400 = $1,510,800

Part B

Cost per Unit = Material + Labor + Variable Overhead + Fixed Overhead.

Last year cost per unit = $1.68 + 1.37 + 1.26 + 3.81= $8.12

This year cost per unit = $2.02 + 1.42 + 1.26 + 5.24 = $9.94

If you have any doubts please comment on the answer.